# In search of the electron-phonon contribution to total energy

### v1

@@ VII Allen’s electron-phonon contribution to the total energy
In Ref. Allen2022, Allen proposed that the contribution from the electron-phonon interaction to the total energy could be computed from the same ingredients as for the zero-point renormalization of the electronic eigenenergies due to the electron-phonon interaction. Using the notation of the present paper, Eq. (4) of Ref. Allen2022 becomes
@@
In a subsequent work [RamasimhaVarma2023], Varma et al. computed the right-hand side of Eq. (113), and argued that it approximates well Eq. (109), the electron-phonon contribution to the total energy. Like Allen, they did not mention that both quantities are equal. We also remark that the eigenvalue renormalization in Eq. (110) is expressed using the adiabatic approximation, but the nonadiabatic version of $\Delta\varepsilon_{n{\mathbf{k}}}$ should be used for infra-red active materials [Ponce2015, Miglio2020].
Let us proceed with the scaling analysis of $E^{\textrm{Allen}}$. Relying on the analysis presented in Sec. III, in particular the fact that the fluctuations of nuclear displacements are proportional to $\lambda$, one finds that $\Delta\varepsilon_{n{\mathbf{k}}}$ scales as $\lambda^{2}$. Actually, $\Delta V_{\nu{\mathbf{q}}}$ scales as $\lambda$, and is present twice in Eq. (110), while $\Delta_{2}V_{\nu{\mathbf{q}}}$ scales like $\lambda^{2}$, see Eqs. (34) and (40) of Ref. Giustino2017. Hence, $E^{\textrm{Allen}}$ also scales as $\lambda^{2}$. It is of second order in the perturbation, not of fourth order. However, the perturbation approach of the previous sections shows that the second-order term gives the phonon contribution to the total energy, no second-order contribution in $\lambda$ can be called an electron-phonon contribution. As we will show, this alleged electron-phonon contribution to the total energy is actually a part of the phonon contribution to the total energy, previously mentioned, $E^{\rm ph}$. Hence $E^{\textrm{Allen}}$ should not be added to $E^{\rm BO}_{|\mathbf{0}}$ and $E^{\rm ph}$. The correct expressions for the lowest-order electron-phonon contribution to the total energy, if any, $E^{\rm elph}$, are given by Eqs. (102a), (106), and (107).
To prove that $E^{\textrm{Allen}}$ is a contribution to $E^{\rm ph}$ and not to $\lambda_{0}^{4}\mathcal{E}^{(4)}$, we further start from Eqs. (1) and (2) of Ref.  Ponce2015:
@@
where the dependence of the electronic energy $E^{\rm e}$ on $\phi$ and $\{{\mathbf{R}}\}$ is omitted for conciseness. In addition, the second-order derivative of the BO total energy is the interatomic force constant, see Eq. (139), which extends for the periodic case to
@@
Therefore, Allen’s energy defined in Ref. Allen2022 is not an electron-phonon energy but a part of the phonon contribution to the total energy. We can inquire about the origin of the problem in Ref. Allen2022. Apparently, the issue stems from the use of an effective Hamiltonian, with separate electronic and phononic terms fixed a priori, to which an electron-phonon interaction is added, changing the total energy. As derived in the present work, following the BO approach, the first-principles BO-DFT phonons come from the BO curvature, which already includes the electronic energy curvature, in addition to nuclear-nuclear energy curvature, and Hartree and exchange-correlation corrections. Hence, phonons cannot be separated from electrons a priori, unlike in the effective Hamiltonian of Ref. Allen2022.
@@ VIII Connection with the Allen-Heine-Cardona theory
Despite $E^{\textrm{Allen}}$ not being the desired electron-phonon interaction energy, one can nonetheless study it and examine its connection with the Allen-Heine-Cardona (AHC) theory of zero-point renormalization. Using Ref. Ponce2015, $\Delta\varepsilon_{n{\mathbf{k}}}$ in Eq. (113) can be decomposed as:
@@
where $\Delta\varepsilon_{n{\mathbf{k}}}^{\rm AHC}$ is obtained from the AHC theory [Allen1976, Allen1981] with the rigid-ion approximation [Ponce2014a] to describe the Debye-Waller term. $\Delta\varepsilon_{n{\mathbf{k}}}^{\rm AHC}$ has the following sum-over-state expression [Gonze2011, Ponce2015]
@@
where $\eta$ should be a positive infinitesimal in the AHC theory. Beyond AHC, $\eta$ would effectively describe the lifetime of the quasiparticle state [Lihm2024]. In this work, without computing $\eta$, it is chosen to be 5 meV.
The non-rigid-ion approximation (NRIA) $\Delta\varepsilon_{n{\mathbf{k}}}^{\rm NRIA}$ terms in Eq. (120) have been shown to be small in solids [Ponce2014a, Ponce2025a]. Computing them requires the second-order derivative of the Hamiltonian, not available with current DFPT implementation, that reads:
@@ IX Validation for diamond at the phonon zone-center
In this section, we compute different quantities mentioned in Secs.VI-VIII using diamond as an example, and compare results from perturbation theory with those from finite differences (FD - frozen phonon approach). Although numbers will first be obtained for a single wavevector at the zone center, the procedure is defined for a grid of wavevectors compatible with a supercell of the primitive cell. Let $N$ be the size ratio between the primitive cell and the supercell, the latter being commensurate with the former, and consider similarly that there are $N$ allowed $\mathbf{q}$} wavevectors that match periodic boundary conditions in the supercell. We define $E^{\rm BO,sup}$, the BO energy of the supercell, that will be evaluated with atoms collectively displaced from their equilibrium position. Precisely, the atom labeled $\kappa$ in the cell $l$, with equilibrium position $\mathbf{R}_{l\kappa}^{0}$, will be displaced by $h$ times the eigendisplacement $\boldsymbol{U}_{\nu\kappa{\mathbf{q}}}$, modulated by the phase $e^{-i{\mathbf{q}}\cdot\mathbf{R}_{l}}$, giving the displaced position
@@
We refer the reader to Ref. Ponce2014, Eqs. (70) and (71), for the treatment of complex displacements, using linear combinations of such displacements. The phonon frequencies are obtained using the following FD formulation:
@@
In this section, we consider the $\mathbf{q}$$\Gamma$ case so that finite difference calculations can be performed only in the primitive cell. We will use the finite $\mathbf{q}$ formulation in Sec. XI. Calculations are performed with the Abinit v10.4.5 software [Gonze2016, Gonze2020, Verstraete2025] with a 40 Ha plane-wave energy cutoff and a zone-centered $\times8$$\times$8 $\mathbf{k}$}-point grid. We rely on the Generalized Gradient Approximation (GGA) PBE [Perdew1996] exchange-correlation functional, with the carbon scalar-relativistic standard PBE pseudopotential from PseudoDojo [vanSetten2018] v0.5. We compute a relaxed 6.7513 Bohr lattice parameter, slightly overestimating the 6.74 Bohr experimental value [Surratt1973]. Diamond is a non-polar materials with a long-range quadrupole value of 2.46 eeBohr [Ponce2021]. We note that the calculation of the quadrupole tensor in solids from perturbation theory has so far only been implemented in the Abinit software [Royo2019, Gonze2020]. Since Abinit v10.4.5, the calculations of quadrupole tensor supports nonlinear core correction (NLCC) of pseudopotentials and has been used here. We obtain a quadrupole value of 2.502 eeBohr using a $\times16$$\times$16 $\mathbf{k}$}-point grid.
@@
This recent development is crucial because the use of a model core charge has been shown to be important [Louie1982] in the case of generalized gradient approximation pseudopotentials such as PBE. In fact, when the valence and core densities overlap, neglecting the core density results in an artificially large gradient of the density. We have numerically tested this effect for diamond and found that ignoring the NLCC leads to numerically unstable finite-difference calculations but is acceptable for perturbation theory. Unless specified otherwise, we use the above-mentioned pseudopotential which includes NLCC.
@@
Table 2: Comparison for diamond between the finite difference (FD) and density functional perturbation theory (DFPT) for different quantities at $\mathbf{q}$$\boldsymbol{\Gamma}$ and T=0 K, at the optimized BO lattice parameters. The $\Delta\varepsilon_{\boldsymbol{\Gamma}}$ refers to the band zero-point renormalization of the triply degenerate valence band maximum, $\Delta\varepsilon_{\boldsymbol{\Gamma}}^{\rm NRIA}$ is the non-rigid ion approximation (NRIA), and $\Delta\varepsilon_{\boldsymbol{\Gamma}}^{\rm AHC}$ refers to the Allen-Heine-Cardona (AHC) theory using the rigid-ion approximation. The 0th-order Born-Oppenheimer (BO) total energy is given by $E^{\rm BO}$ and include the electronic part is given by $E^{\rm e}$, and the other terms $E^{\rm other}$. The second-order contribution to total energy due to phonons is $E^{\rm ph}$. It can be decomposed into half the electronic contribution $E^{\rm Allen}$, and the rest $E^{\rm ph,other}$. The electron-phonon part of the fourth-order contribution to total energy $E^{\rm elph}$ can be computed using DFPT. For total energies, we use a $\times8$$\times$8 $\mathbf{k}$-point grid with symmetry, detailed in Table 3. Absolute total energy values depend on the choice of pseudopotential but relative energy differences are observables.
@@
$\omega_{\boldsymbol{\Gamma}}$ DFPT (139) 159.989
@@
We start by verifying phonon frequencies computed by perturbation theory and finite difference (FD) in Table 2 using a $\times8$$\times$8 $\mathbf{k}$}-grid, since phonon frequencies will be used in later quantities. All finite-difference calculations are performed using a Richardson extrapolation with Romberg order 4, starting from hh=2. We find that the triply degenerate optical modes agree within 0.013 meV between the density functional perturbation theory (DFPT) [Gonze1997, Baroni2001] and FD values. We then compute the phonon zone-center contribution to the zero-point renormalization (ZPR) of the valence band maximum (VBM) computed using FD and obtain 27.695 meV, close to the value of 28.428 meV of a prior study [Ponce2014a]. Since the VBM is triply degenerate in band indices and any linear combination is a valid one, we report the average value over the three bands. The same value cannot be directly computed using DFPT as it requires a higher-order perturbation currently not available in first-principles software. However, it can be approximated using the RIA which gives 23.806 meV, again in close agreement with the 24.830 meV value of a previous study [Ponce2014a]. The neglected NRIA term can be computed using FD, and then added to the RIA term giving 27.700 meV, in very close agreement with the direct FD value. We note that the RIA is a good approximation for solids when integrated over $\mathbf{q}$-points [Ponce2014a, Ponce2025a]. As a technical note, it is possible in Abinit to reduce the computation cost by using a reduced ratio between the radius of the $\mathbf{G}$-vector sphere that can be inserted in the real-space fast Fourier transform box and the $\mathbf{G}$-vector sphere used to represent the wavefunction. However, we find here that it is crucial to use a ratio of 2 to have an exact density, see Table 6 of the Appendix D.
@@
Next, we report in Table 2 the zeroth-order BO total energy. Such total energy, provided as an output of a standard self-consistent DFT cycle, contains a core electron energy, that is assumed constant, and ignored via the use of pseudopotential. Therefore the total energy depends on the arbitrary choice of the pseudopotential. In contrast, the contribution from its eigenenergy component $E^{\rm e}$ is only weakly dependent on the choice of pseudopotential, in a way similar to other predicted observables with DFT. We compute a value of 10.119 eV directly from the output of the Abinit software and verify that it gives the same result as the direct sum over the occupied eigenvalues $\sum_{n{\mathbf{k}}}\varepsilon_{n{\mathbf{k}}}f_{n{\mathbf{k}}}$. We then compute the second-order contribution to the total energy due to phonons and find a value of 240.040 meV using DFPT, which is less than 0.04 meV larger than the two FD results. We see that such a contribution is significantly lower than the $E^{\rm e}$ energy contribution.
Table 3: Comparison of the sum over all occupied bands (1-4) of the ZPR, between the finite difference (FD) and Allen-Heine-Cardona (AHC) theory, for diamond with 40 Ha cutoff. The contribution of each k-point is listed, with the corresponding weight. $\times8$$\times$8 $\mathbf{k}$}-grid, standard scalar-relativistic PBE pseudopotential from PseudoDojo. For the FD, Richardson extrapolation with Romberg order 4, starting from a $h=2$ displacement.
$\mathbf{k}$-point weight $\Delta\varepsilon_{1-4,\mathbf{k}}$ $\Delta\varepsilon_{1-4,\mathbf{k}}^{\rm NRIA}$ Diff. $\Delta\varepsilon_{1-4,\mathbf{k}}^{\rm AHC}$
(meV) (meV) (meV) (meV)
0, 0, 0 $\frac{1}{512}$ 139.369 25.441 113.928 113.952
$\frac{1}{8},0,0$ $\frac{8}{512}$ -76.081 21.885 -97.966 -98.186
$\frac{1}{4},0,0$ $\frac{8}{512}$ -78.347 19.149 -97.496 -97.762
$\frac{3}{8},0,0$ $\frac{8}{512}$ -39.933 18.013 -57.946 -58.217
$\frac{1}{2},0,0$ $\frac{4}{512}$ -24.155 17.710 -41.865 -42.136
$\frac{1}{8},\frac{1}{8},0$ $\frac{6}{512}$ -100.133 21.769 -121.902 -121.681
$\frac{1}{4},\frac{1}{8},0$ $\frac{24}{512}$ -100.774 18.400 -119.174 -120.007
$\frac{3}{8},\frac{1}{8},0$ $\frac{24}{512}$ -78.439 16.929 -95.368 -96.023
$\frac{1}{2},\frac{1}{8},0$ $\frac{24}{512}$ -64.532 16.769 -81.301 -81.435
-$\frac{3}{8},\frac{1}{8},0$ $\frac{24}{512}$ -67.085 17.484 -84.569 -84.113
-$\frac{1}{4},\frac{1}{8},0$ $\frac{24}{512}$ -83.279 19.005 -102.284 -101.285
-$\frac{1}{8},\frac{1}{8},0$ $\frac{12}{512}$ -93.054 21.372 -114.426 -113.031
$\frac{1}{4},\frac{1}{4},0$ $\frac{6}{512}$ -97.988 19.000 -116.988 -116.658
$\frac{3}{8},\frac{1}{4},0$ $\frac{24}{512}$ -88.777 16.678 -105.455 -105.796
$\frac{1}{2},\frac{1}{4},0$ $\frac{24}{512}$ -84.746 15.773 -100.519 -100.471
-$\frac{3}{8},\frac{1}{4},0$ $\frac{24}{512}$ -85.456 16.128 -101.584 -100.957
-$\frac{1}{4},\frac{1}{4},0$ $\frac{12}{512}$ -89.319 17.483 -106.802 -105.630
$\frac{3}{8},\frac{3}{8},0$ $\frac{6}{512}$ -78.257 17.372 -95.629 -95.258
$\frac{1}{2},\frac{3}{8},0$ $\frac{24}{512}$ -80.215 15.980 -96.195 -95.960
-$\frac{3}{8},\frac{3}{8},0$ $\frac{12}{512}$ -84.827 15.692 -100.519 -99.836
$\frac{1}{2},\frac{1}{2},0$ $\frac{3}{512}$ -72.174 16.750 -88.924 -88.543
$\frac{3}{8},\frac{1}{4},\frac{1}{8}$ $\frac{24}{512}$ -95.325 19.478 -114.803 -112.791
$\frac{1}{2},\frac{1}{4},\frac{1}{8}$ $\frac{48}{512}$ -89.036 17.192 -106.228 -104.996
-$\frac{3}{8},\frac{1}{4},\frac{1}{8}$ $\frac{24}{512}$ -85.732 16.339 -102.071 -101.185
$\frac{1}{2},\frac{3}{8},\frac{1}{8}$ $\frac{24}{512}$ -88.837 16.676 -105.513 -104.352
-$\frac{3}{8},\frac{3}{8},\frac{1}{8}$ $\frac{48}{512}$ -90.944 15.234 -106.178 -105.401
-$\frac{1}{4},\frac{3}{8},\frac{1}{8}$ $\frac{24}{512}$ -92.176 15.392 -107.568 -106.800
-$\frac{3}{8},\frac{1}{2},\frac{1}{8}$ $\frac{12}{512}$ -86.008 15.087 -101.095 -100.799
-$\frac{1}{4},\frac{1}{2},\frac{1}{4}$ $\frac{6}{512}$ -94.059 13.946 -108.005 -107.826
EAllen EAllen,NRIA Diff. EAllen,AHC
k-integrated -84.214 17.086 -101.300 -100.783
@@
To assess the eigenenergy contribution to that second-order energy, we compute the ZPR for all $\mathbf{k}$-point of an homogeneous $\times8$$\times$8 $\mathbf{k}$}-grid using crystal symmetry, and report their values in Table 3. We find a close agreement between perturbation theory and FD, with less than 1.5 meV difference. From its weighted summation, we find that the eigenenergy contribution to the second-order energy is -84.214 meV, which amounts to over one-third the total second-order contribution. This value can be approximated via DFPT and the RIA which gives 100.783-100.783 meV, overestimating by 20% the true value. As mentioned earlier, we expect the RIA to improve when summing over dense $\mathbf{q}$-points grids.
@@
Figure 3: Diamond fourth-order total energy contribution computed using DFPT, Eq. (106), and with a sum-over-state (SOS) expression, Eq. (107). The $\mathbf{k}$-point grid used is $\times8$$\times$8 with only a single $\mathbf{q}=\boldsymbol{\Gamma}$ point.
@@
Finally, we compute the electron-phonon contribution to the fourth-order total energy using DFPT, $E^{\rm elph}$, Eq. (106), and obtain 3.639 meV. Alternatively, we can use Eq. (107) to compute the same quantity via the electron-phonon interaction matrix elements. As shown in Fig. 3, both result in the same value but Eq. (107) requires a sum over empty states which converges slowly. In the case of diamond, we need at least 300 bands to converge that formulation. Such a small value is nearly two orders of magnitude smaller than the second-order contribution, which is expected given the $M_{0}^{-1/2}$ scaling between each higher-even order terms. Although small, this contribution could play a role in stabilizing an allotrope over another, especially in the case of magnetic materials that can have a very shallow energy landscape.
X Size consistency

### v2

@@ VII Allen’s electron-phonon contribution to the total energy
In Ref. 6, Allen proposed that the contribution from the electron-phonon interaction to the total energy could be computed from the same ingredients as for the zero-point renormalization of the electronic eigenenergies due to the electron-phonon interaction. Using the notation of the present paper, Eq. (4) of Ref. 6 becomes
@@
In a subsequent work [68], Varma et al. computed the right-hand side of Eq. (109), and argued that it approximates well Eq. (105), the electron-phonon contribution to the total energy. Like Allen, they did not mention that both quantities are equal. We also remark that the eigenvalue renormalization in Eq. (106) is expressed using the adiabatic approximation, but the nonadiabatic version of $\Delta\varepsilon_{n{\mathbf{k}}}$ should be used for infra-red active materials [60, 52].
Let us proceed with the mass-scaling analysis of $E^{\textrm{Allen}}$. Relying on the analysis presented in Sec. III, in particular the fact that the fluctuations of nuclear displacements are proportional to $\lambda$, one finds that $\Delta\varepsilon_{n{\mathbf{k}}}$ scales as $\lambda^{2}$. Actually, $\Delta V_{\nu{\mathbf{q}}}$ scales as $\lambda$, and is present twice in Eq. (106), while $\Delta_{2}V_{\nu{\mathbf{q}}}$ scales like $\lambda^{2}$, see Eqs. (34) and (40) of Ref. 26. Hence, $E^{\textrm{Allen}}$ also scales as $\lambda^{2}$. It is of second order in the perturbation, not of fourth order. However, the perturbation approach of the previous sections shows that the second-order term gives the phonon contribution to the total energy, no second-order contribution in $\lambda$ can be called an electron-phonon contribution. As we will show, this alleged electron-phonon contribution to the total energy is actually a part of the phonon contribution to the total energy, previously mentioned, $E^{\rm ph}$. Hence $E^{\textrm{Allen}}$ should not be added to $E^{\rm BO}_{|\mathbf{0}}$ and $E^{\rm ph}$. The correct expressions for the lowest-order electron-phonon contribution to the total energy, if any, $E^{\rm elph}$, are given by Eqs. (98a), (102), and (103).
To prove that $E^{\textrm{Allen}}$ is a contribution to $E^{\rm ph}$ and not to $\lambda_{0}^{4}\mathcal{E}^{(4)}$, we further start from Eqs. (1) and (2) of Ref.  60:
@@
where the dependence of the electronic energy $E^{\rm e}$ on $\phi$ and $\{{\mathbf{R}}\}$ is omitted for conciseness. In addition, the second-order derivative of the BO total energy is the interatomic force constant, see Eq. (LABEL:eq:secular), which extends for the periodic case to
@@
Therefore, Allen’s energy defined in Ref. 6 is not an electron-phonon energy but a part of the phonon contribution to the total energy. We can inquire about the origin of the problem in Ref. 6. Apparently, the issue stems from the use of an effective Hamiltonian, with separate electronic and phononic terms fixed a priori, to which an electron-phonon interaction is added, changing the total energy. As derived in the present work, following the BO approach, the first-principles BO-DFT phonons come from the BO curvature, which already includes the electronic energy curvature, in addition to nuclear-nuclear energy curvature, and Hartree and exchange-correlation corrections. Hence, phonons cannot be separated from electrons a priori, unlike in the effective Hamiltonian of Ref. 6.
@@ VIII Connection with the Allen-Heine-Cardona theory
Despite $E^{\textrm{Allen}}$ not being the desired electron-phonon interaction energy, one can nonetheless study it and examine its connection with the Allen-Heine-Cardona (AHC) theory of zero-point renormalization. Using Ref. 60, $\Delta\varepsilon_{n{\mathbf{k}}}$ in Eq. (109) can be decomposed as:
@@
where $\Delta\varepsilon_{n{\mathbf{k}}}^{\rm AHC}$ is obtained from the AHC theory [2, 3] with the rigid-ion approximation [59] to describe the Debye-Waller term. $\Delta\varepsilon_{n{\mathbf{k}}}^{\rm AHC}$ has the following sum-over-state expression [27, 60]
@@
where $\eta$ should be a positive infinitesimal in the AHC theory. Beyond AHC, $\eta$ would effectively describe the lifetime of the quasiparticle state [44]. In this work, without computing $\eta$, it is chosen to be 5 meV.
The non-rigid-ion approximation (NRIA) $\Delta\varepsilon_{n{\mathbf{k}}}^{\rm NRIA}$ terms in Eq. (116) have been shown to be small in solids [59, 63]. Computing them requires the second-order derivative of the Hamiltonian, not available with current DFPT implementation, that reads:
@@ IX Validation for diamond at the phonon zone-center
In this section, we compute different quantities mentioned in Secs.VI-VIII using diamond as an example, and compare results from perturbation theory with those from finite differences (FD - frozen phonon approach). Although the numbers will first be obtained for a single wavevector at the zone center, the procedure is defined for a grid of wavevectors compatible with a supercell of the primitive cell.
Let $N$ be the size ratio between the primitive cell and the supercell, the latter being commensurate with the former, and consider similarly that there are $N$ allowed $\mathbf{q}$} wavevectors that match periodic boundary conditions in the supercell. We define $E^{\rm BO,sup}$, the BO energy of the supercell, that will be evaluated with atoms collectively displaced from their equilibrium position. Precisely, the atom labeled $\kappa$ in the cell $l$, with equilibrium position $\mathbf{R}_{l\kappa}^{0}$, will be displaced by $h$ times the eigendisplacement $\boldsymbol{U}_{\nu\kappa{\mathbf{q}}}$, modulated by the phase $e^{-i{\mathbf{q}}\cdot\mathbf{R}_{l}}$, giving the displaced position
@@
We refer the reader to Ref. 58, Eqs. (70) and (71), for the treatment of complex displacements, using linear combinations of such displacements. The phonon frequencies are obtained using the following FD formulation:
@@
In this section, we consider the $\mathbf{q}$$\Gamma$ case so that finite difference calculations can be performed only in the primitive cell. We will use the finite $\mathbf{q}$ formulation in Sec. X. Calculations are performed with the Abinit v10.4.5 software [28, 29, 90] with a 40 Ha plane-wave energy cutoff and a zone-centered $\times8$$\times$8 $\mathbf{k}$}-point grid. We rely on the Generalized Gradient Approximation (GGA) PBE [57] exchange-correlation functional, with the carbon scalar-relativistic standard PBE pseudopotential from PseudoDojo [87] v0.5. We compute a relaxed 6.7513 Bohr lattice parameter, slightly overestimating the 6.74 Bohr experimental value [85]. Diamond is a non-polar material with a long-range quadrupole value of 2.46 eeBohr [64]. We note that the calculation of the quadrupole tensor in solids from perturbation theory has so far only been implemented in the Abinit software [75, 29]. Since Abinit v10.4.5, the calculations of quadrupole tensor supports nonlinear core correction (NLCC) of pseudopotentials and has been used here. We obtain a quadrupole value of 2.502 eeBohr using a $\times16$$\times$16 $\mathbf{k}$}-point grid.
@@
This recent development is crucial because the use of a model core charge has been shown to be important [47] in the case of generalized gradient approximation pseudopotentials such as PBE. In fact, when the valence and core densities overlap, neglecting the core density results in an artificially large gradient of the density. We have numerically tested this effect for diamond and found that ignoring the NLCC leads to numerically unstable finite-difference calculations but is acceptable for perturbation theory. Unless specified otherwise, we use the above-mentioned pseudopotential which includes NLCC.
@@
Table 2: Comparison for diamond between the finite difference (FD) and density functional perturbation theory (DFPT) for different quantities at $\mathbf{q}$$\boldsymbol{\Gamma}$ and T=0 K, at the optimized BO lattice parameters. The $\Delta\varepsilon_{\boldsymbol{\Gamma}}$ refers to the band zero-point renormalization of the triply degenerate valence band maximum, $\Delta\varepsilon_{\boldsymbol{\Gamma}}^{\rm NRIA}$ is the non-rigid ion approximation (NRIA), and $\Delta\varepsilon_{\boldsymbol{\Gamma}}^{\rm AHC}$ refers to the Allen-Heine-Cardona (AHC) theory using the rigid-ion approximation. The 0th-order Born-Oppenheimer (BO) total energy is given by $E^{\rm BO}$ and include the electronic part is given by $E^{\rm e}$, and the other terms $E^{\rm other}$. The second-order contribution to total energy due to phonons is $E^{\rm ph}$. It can be decomposed into half the electronic contribution $E^{\rm Allen}$, and the rest $E^{\rm ph,other}$. The electron-phonon part of the fourth-order contribution to total energy $E^{\rm elph}$ can be computed using DFPT. Energies are given per unit cell (2 atoms). For total energies, we use a $\times8$$\times$8 $\mathbf{k}$-point grid with symmetry, detailed in Table LABEL:tab:table2. Absolute total energy values depend on the choice of pseudopotential but relative energy differences are observables.
@@
$\omega_{\boldsymbol{\Gamma}}$ DFPT (LABEL:eq:secular) 159.989
@@
We start by verifying phonon frequencies computed by perturbation theory and finite difference (FD) in Table 2 using a $\times8$$\times$8 $\mathbf{k}$}-grid, since phonon frequencies will be used in later quantities. All finite-difference calculations are performed using a Richardson extrapolation with Romberg order 4, starting from hh=2. We find that the triply degenerate optical modes agree within 0.013 meV between the density functional perturbation theory (DFPT) [33, 8] and FD values. We then compute the phonon zone-center contribution to the zero-point renormalization (ZPR) of the valence band maximum (VBM) computed using FD and obtain 27.695 meV, close to the value of 28.428 meV of a prior study [59]. Since the VBM is triply degenerate in band indices and any linear combination is a valid one, we report the average value over the three bands. The same value cannot be directly computed using DFPT as it requires a higher-order perturbation currently not available in first-principles software. However, it can be approximated using the RIA which gives 23.806 meV, again in close agreement with the 24.830 meV value of a previous study [59]. The neglected NRIA term can be computed using FD, and then added to the RIA term giving 27.700 meV, in very close agreement with the direct FD value. We note that the RIA is a good approximation for solids when integrated over $\mathbf{q}$-points [59, 63]. As a technical note, it is possible in Abinit to reduce the computation cost by using a reduced ratio between the radius of the $\mathbf{G}$-vector sphere that can be inserted in the real-space fast Fourier transform box and the $\mathbf{G}$-vector sphere used to represent the wavefunction. However, we find here that it is crucial to use a ratio of 2 to have an exact density, see Sec. LABEL:app:improvement of the SI.
@@
Next, we report in Table 2 the zeroth-order BO total energy. Such total energy, provided as an output of a standard self-consistent DFT cycle, contains a core electron energy, that is assumed constant, and ignored via the use of pseudopotential. Therefore, the total energy depends on the arbitrary choice of the pseudopotential. In contrast, the contribution from its eigenenergy component $E^{\rm e}$ is only weakly dependent on the choice of pseudopotential, in a way similar to other predicted observables with DFT. We compute a value of 10.119 eV directly from the output of the Abinit software and verify that it gives the same result as the direct sum over the occupied eigenvalues $\sum_{n{\mathbf{k}}}\varepsilon_{n{\mathbf{k}}}f_{n{\mathbf{k}}}$. We then compute the second-order contribution to the total energy due to phonons and find a value of 240.040 meV using DFPT, which is less than 0.04 meV larger than the two FD results. We see that such a contribution is significantly lower than the $E^{\rm e}$ energy contribution.
@@
To assess the eigenenergy contribution to that second-order energy, we compute the ZPR for all $\mathbf{k}$-point of an homogeneous $\times8$$\times$8 $\mathbf{k}$}-grid using crystal symmetry, and report their values in Sec. LABEL:sec:zpr of the SI. We find a close agreement between perturbation theory and FD, with less than 1.5 meV difference. From its weighted summation, we find that the eigenenergy contribution to the second-order energy is -84.214 meV, which amounts to over one-third the total second-order contribution. This value can be approximated via DFPT and the RIA which gives 100.783-100.783 meV, overestimating by 20% the true value. As mentioned earlier, we expect the RIA to improve when summing over dense $\mathbf{q}$-points grids.
@@
Figure 3: Diamond fourth-order total energy contribution (per unit cell) computed using DFPT, Eq. (102), and with a sum-over-state (SOS) expression, Eq. (103). The $\mathbf{k}$-point grid used is $\times8$$\times$8 with only a single $\mathbf{q}=\boldsymbol{\Gamma}$ point.
@@
Furthermore, we compute the electron-phonon contribution to the fourth-order total energy using DFPT, $E^{\rm elph}$, Eq. (102), and obtain 3.639 meV. Alternatively, we can use Eq. (103) to compute the same quantity via the electron-phonon interaction matrix elements. As shown in Fig. 3, both result in the same value but Eq. (103) requires a sum over empty states which converges slowly. In the case of diamond, we need at least 300 bands to converge that formulation. Such a small value is nearly two orders of magnitude smaller than the second-order contribution, which is expected given the $M_{0}^{-1/2}$ scaling between each higher-even order terms. Although small, this contribution could play a role in stabilizing an allotrope over another, especially in the case of magnetic materials that can have a very shallow energy landscape.